5. Compute the equivalent units of production for conversion.

6. What is the cost of beginning work in process inventory plus the cost added during the period for materials?

7. What is the cost of beginning work in process inventory plus the cost added during the period for conversion?

8. What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.)

Clopack Company manufactures one product that goes through one processing department called Mixing. All raw materials are introduced at the start of work in the Mixing Department. The company uses the weighted-average method of process costing. Its Work in Process T-account for the Mixing Department for June follows (all forthcoming questions pertain to June):

Work in Process—Mixing Department
June 1 balance 36,000 Completed and transferred to Finished Goods ?
Materials 122,440
Direct labor 81,500
Overhead 99,000
June 30 balance ?

The June 1 work in process inventory consisted of 5,200 units with $19,130 in materials cost and $16,870 in conversion cost. The June 1 work in process inventory was 100% complete with respect to materials and 60% complete with respect to conversion. During June, 37,700 units were started into production. The June 30 work in process inventory consisted of 8,400 units that were 100% complete with respect to materials and 50% complete with respect to conversion.
Answer #1

Unit transferred out = 5200+37700-8400 = 34500

5) equivalent unit of conversion = 34500+(8400*50%) = 38700 Units

6) Cost of beginning work in process plus cost added for material = 19130+122440 = 141570

7) Cost of beginning work in process plus cost added for conversion = 16870+81500+99000 = 197370

8) Cost per equivalent unit of material = 141570/42900 = 3.30